Use standard reduction potentials to calculate the equilibrium constant for the reaction:

I2(s) + Hg(l)2I-(aq) + Hg2+(aq)


Hint: Carry at least 5 significant figures during intermediate calculations to avoid round off error when taking the antilogarithm.

a.) Find the equilibrium constant

b.) G° for this reaction would be greater/less than zero.
Answer #1

Given grun S2 (8) + High) Phant Honda At anode Coundation ) Hg – Hy? #ze, Mga+1+19 =0,85 \V At cathode creduction) Ia + Ré -In keq = -24: 14864 e keq = 18-24. 14864 2.303 Keg en keq = -24.14664 2.303 log keg= -24. 14664 log keg = – 10.485732 i. Keg=

Thanks!